# Formation of GW230529 from Isolated Binary Evolution and Its Electromagnetic Counterparts

### v2

@@ 3.1 Tidal Disruption Probability
Figure 3 shows the parameter space where the NS can be tidally disrupted using Equation (1). For a given $\chi_{{\rm BH},z}$, BHNS mergers composed of a low-mass BH component and a low-mass NS component are more easily to have tidal disruption. The mass space that allows NS tidal disruption expands significantly with increasing $\chi_{{\rm BH},z}$ and the adoption of a stiffer EoS. We display the 90% credible posterior distributions and the medians of component masses, as well as the medians of BH aligned-spins $\chi_{{\rm BH},z}$, for GW200105, GW200115, and GW230529 (Abbott et al., 2021a; The LIGO Scientific Collaboration et al., 2024) in Figure 3. Based on the GW observed samples reported by the LVK Collaboration (Abbott et al., 2021a; The LIGO Scientific Collaboration et al., 2024), the explicit results of the tidal disruption probabilities for these three BHNS mergers are listed in Table 2. It is obvious that GW200105 is unlikely to make tidal disruption to generate any EM counterparts, because its BH and NS components are massive with the mass distribution significantly outside the tidal disruption region. Although GW200115 has BH and NS masses much lighter than those of GW200105, tidal disruption can occur with only a very low probability of $P_{\rm tidal}=2.76\%$ under the EoS of DD2. Compared with GW200105 and GW200115 whose BH masses are $\gtrsim 5\,M_{\odot}$, GW230529, as a mgBHNS merger, occupies a larger portion of its mass parameter space located within the disruption region, especially when considering a stiffer EoS. More specifically, the tidal disruption probabilities are $12.8\%$ and $63.2\%$ by adopting the EoS of AP4 and DD2, respectively. The former value is similar to that reported in The LIGO Scientific Collaboration et al. (2024), which was obtained by marginalizing over the EoS.
@@ 3.4 Implication of EM detectability for BHNS Mergers
Based on our population synthesis simulation results, we then simulate the tidal disruption probability, kilonova detectable rate, and GRB detection rate for GW BHNS mergers detected in the near future GW observing era (e.g., O4b and O5). We conservatively consider BHNS mergers occurring within $300\,{\rm Mpc}$, which is the observed distance of GW200105 and GW200115, as well as the representative distance of detectable BHNS mergers in O4 (Abbott et al., 2018b; Zhu et al., 2021b; Colombo et al., 2023; Gupta et al., 2023). We find that $100\%$ and $\gtrsim 90\%$ disrupted BHNS mergers are expected to originate from mgBHNS mergers by adopting the EoS of AP4 and DD2, respectively. When the detection depth exceeds $24\,{\rm mag}$, the survey projects can cover the majority of kilonovae associated with BHNS mergers. However, considering the rapid rise and decline of kilonovae, their signals may be challenging to be identified. Additionally, we find that the detection rate of GRBs originating from mgBHNS mergers is $\sim 0.1-0.4\,{\rm yr}^{-1}$. Therefore, in the future, it may be possible to search for kilonova signals via the target-of-opportunity follow-up observations of both GW and GRB triggers, like the multimessenger observations between GW170817/GRB170817A/AT2017gfo.
@@ 4 Conclusion
In this Letter, we use COMPAS to explore the formation of GW230529, which is a mgBHNS merger between a $3.6^{+0.7}_{-1.2}M_{\odot}$ BH and a $1.43^{+0.59}_{-0.19}M_{\odot}$ NS reported by the LVK Collaboration. By adopting the ‘delayed’ SN prescription, our population synthesis simulations can simultaneously match the inferred event rate densities of the mgBHNS and total BHNS mergers obtained from the population analyses, along with the population distribution of the BH mass in BHNS mergers modeled by the LVK Collaboration. The mass posterior of GW230529 significantly overlaps with the highest probability region of the simulated BHNS population. Furthermore, mgBHNS mergers can also originate from dynamical formation in dense stellar environments or triple system, but the expected rate densities could be much lower than inferred GW mgBHNS rate densities (e.g., Fragione et al., 2020; Gupta et al., 2020; Ye et al., 2020; Tagawa et al., 2021). Since GW230529 contributes significantly to the current inferred mgBHNS rate densities, we suggest that GW230529 could originate from the isolated binary evolution channel.
@@
Population Synthesis Simulations in this Letter made use of the COMPAS rapid binary population synthesis code (version 02.39.00), which is freely available at http://github.com/TeamCOMPAS/COMPAS. We thank the GWOSC (Gravitational Wave Open Science Center; https://gwosc.org) for making the data of GW230529 publicly available at https://doi.org/10.7935/6k89-7q62. We thank Floor Broekgaarden for publicly providing access to a template of a table that can summarize the binary population synthesis settings (https://github.com/FloorBroekgaarden/templateForTableBPSsettings). We thank Floor Broekgaarden, Paul D. Lasky, Liang-Duan Liu, Ilya Mandel, Eric Thrane, Xiangyu Ivy Wang, Zepei Xing, and Yun-Wei Yu for helpful discussions. JPZ thanks the COMPAS group at Monash University. YK and LS were supported by the Beijing Natural Science Foundation (1242018), the National SKA Program of China (2020SKA0120300), and the Max Planck Partner Group Program funded by the Max Planck Society. YQ acknowledges the support of Anhui Provincial Natural Science Foundation (grant No. 2308085MA29). We summarize the binary population synthesis settings of our COMPAS simulations in Table 4.

### v3

@@ 3.1 Tidal Disruption Probability
Figure 3 shows the parameter space where the NS can be tidally disrupted using Equation (1). For a given $\chi_{{\rm BH},z}$, BHNS mergers composed of a low-mass BH component and a low-mass NS component are more easily to have tidal disruption. The mass space that allows NS tidal disruption expands significantly with increasing $\chi_{{\rm BH},z}$ and the adoption of a stiffer EoS. We display the 90% credible posterior distributions and the medians of component masses, as well as the medians of BH aligned-spins $\chi_{{\rm BH},z}$, for GW200105, GW200115, and GW230529 (Abbott et al., 2021a; The LIGO Scientific Collaboration et al., 2024) in Figure 3. Based on the observed GW samples reported by the LVK Collaboration (Abbott et al., 2021a; The LIGO Scientific Collaboration et al., 2024), the explicit results of the tidal disruption probabilities for these three BHNS mergers are listed in Table 2. It is obvious that GW200105 is unlikely to make tidal disruption to generate any EM counterparts, because its BH and NS components are massive with the mass distribution significantly outside the tidal disruption region. Although GW200115 has BH and NS masses much lighter than those of GW200105, tidal disruption can occur with only a very low probability of $P_{\rm tidal}=2.76\%$ under the EoS of DD2. Compared with GW200105 and GW200115 whose BH masses are $\gtrsim 5\,M_{\odot}$, GW230529, as a mgBHNS merger, occupies a larger portion of its mass parameter space located within the disruption region, especially when considering a stiffer EoS. More specifically, the tidal disruption probabilities are $12.8\%$ and $63.2\%$ by adopting the EoS of AP4 and DD2, respectively. The former value is similar to that reported in The LIGO Scientific Collaboration et al. (2024), which was obtained by marginalizing over the EoS.
@@ 3.4 Implication of EM detectability for BHNS Mergers
Based on our $\alpha 2\sigma 265$ population synthesis simulation result, we then simulate the tidal disruption probability, kilonova detectable rate, and GRB detection rate for GW BHNS mergers detected in the near future GW observing era (e.g., O4b and O5). We conservatively consider BHNS mergers occurring within $300\,{\rm Mpc}$, which is the observed distance of GW200105 and GW200115, as well as the representative distance of detectable BHNS mergers in O4 (Abbott et al., 2018b; Zhu et al., 2021b; Colombo et al., 2023; Gupta et al., 2023). The results are listed in Table 3 without considering the simulated errors. We find that $100\%$ and $\gtrsim 90\%$ disrupted BHNS mergers are expected to originate from mgBHNS mergers by adopting the EoS of AP4 and DD2, respectively. When the detection depth exceeds $24\,{\rm mag}$, the survey projects can cover the majority of kilonovae associated with BHNS mergers. However, considering the rapid rise and decline of kilonovae, their signals may be challenging to be identified. Additionally, we find that the detection rate of GRBs originating from mgBHNS mergers is $\sim 0.1-0.4\,{\rm yr}^{-1}$. Therefore, in the future, it may be possible to search for kilonova signals via the target-of-opportunity follow-up observations of both GW and GRB triggers, like the multimessenger observations between GW170817/GRB170817A/AT2017gfo.
@@ 4 Conclusion
In this paper, we use COMPAS to explore the formation of GW230529, which is a mgBHNS merger between a $3.6^{+0.7}_{-1.2}M_{\odot}$ BH and a $1.43^{+0.59}_{-0.19}M_{\odot}$ NS reported by the LVK Collaboration. By adopting the ‘delayed’ SN prescription, our population synthesis simulations can simultaneously match the inferred event rate densities of the mgBHNS and total BHNS mergers obtained from the population analyses, along with the population distribution of the BH mass in BHNS mergers modeled by the LVK Collaboration. The mass posterior of GW230529 significantly overlaps with the highest probability region of the simulated BHNS population. Furthermore, mgBHNS mergers can also originate from dynamical formation in dense stellar environments or triple system, but the expected rate densities could be much lower than inferred GW mgBHNS rate densities (e.g., Fragione et al., 2020; Gupta et al., 2020; Ye et al., 2020; Tagawa et al., 2021). Since GW230529 contributes significantly to the current inferred mgBHNS rate densities, we suggest that GW230529 could originate from the isolated binary evolution channel.
@@
Population Synthesis Simulations in this paper made use of the COMPAS rapid binary population synthesis code (version 02.39.00), which is freely available at http://github.com/TeamCOMPAS/COMPAS. We thank the GWOSC (Gravitational Wave Open Science Center; https://gwosc.org) for making the data of GW230529 publicly available at https://doi.org/10.7935/6k89-7q62. We thank Floor Broekgaarden for publicly providing access to a template of a table that can summarize the binary population synthesis settings (https://github.com/FloorBroekgaarden/templateForTableBPSsettings). The authors acknowledge anonymous referees for useful discussions. We thank Floor Broekgaarden, Paul D. Lasky, Liang-Duan Liu, Ilya Mandel, Eric Thrane, Xiangyu Ivy Wang, Zepei Xing, and Yun-Wei Yu for helpful discussions. JPZ thanks the COMPAS group at Monash University. JPZ acknowledges support from the Australian Research Council Centre of Excellence for Gravitational Wave Discovery (OzGrav), through project number CE17010004. YK and LS were supported by the Beijing Natural Science Foundation (1242018), the National SKA Program of China (2020SKA0120300), and the Max Planck Partner Group Program funded by the Max Planck Society. YQ acknowledges the support of Anhui Provincial Natural Science Foundation (grant No. 2308085MA29). We summarize the binary population synthesis settings of our COMPAS simulations in Table 4.